The length of a rectangular carpet is 5.5' greater than twice the width. The area is 90. Find the dimensions
Answer by mananth(16946)

let the width be x feet
length = 2x+5.5
Area = 90
..
x(2x+11/2)=90
x(4x+11 /2 ) =90
4x^2+11x/2 =90
4x^2+11x=180
4x^2+11x-180=0
x1, x2 the roots of the equation are
x1=(-11+sqrt((11^2)+4*4*180)) /( 2*8)
x1=5.47 feet the width
..
x2=(-11-sqrt((11^2)+4*4*180)) /( 2*8)
x2=-8.22 (ignore)
Length = 2x+5.5
=2*5.47+5.5
=16.44 feet
..
CHECK
5.47*16.44=89.92